OpenStudy (anonymous):

You are interested in purchasing a wide-screen television set at Target. On this type of TV, the ratio of screen width to screen height is 16 to 9. If a certain model you are considering has a screen width of 48 inches, what is the height of the screen

OpenStudy (anonymous):

What is the height of the screen (in inches)?

OpenStudy (nurali):

16/9 = 48/x Cross Multiply 16x=48*9 16x=432 x=432/16 x=________

OpenStudy (anonymous):

9/16 x 48

OpenStudy (anonymous):

x=432/16..do you divide that or 9/16 x48

OpenStudy (anonymous):

27 inches

OpenStudy (anonymous):

yea that's the same answer I got when I divided 432/16